How many quarts of a 50% solution of acid must be added to 20 quarts of a 20% solution of acid to obtain a mixture containing a 40% solution of acid?
Found 2 solutions by ankor@dixie-net.com, richwmiller:
Answer by ankor@dixie-net.com(22740)

How many quarts of a 50% solution of acid must be added to 20 quarts of a 20% solution of acid to obtain a mixture containing a 40% solution of acid?
:
let x = amt of the 50% solution
:
A typical mixture equation using the decimal equiv
.50x + .20(20) = .40(x+20)
.50x + 4 = .40x + 8
.50x - .40x = 8 - 4
.10x = 4
x = 4/.1
x = 40 qts of 50% solution

Answer by richwmiller(17219)

We have 20 qts of 20% solution.
Using the method of alligation.
20= 40 - 20
10= 50 - 40
40.0= 20/(10/20) qts of 50% solution to add to your 20 qts of 20% solution.
60.0= (20+40.0) qts of the final 40% solution.
check
0.5 * 40.0 + 0.2 * 20 = 0.4*(20+40.0)
20.0 + 4.0 = 0.4*(60.0)
24.0 = 24.0
ok
20         10        20 qts of 20% solution
      40            60 qts of 40% solution
50         20 20/10*20=40 qts of 50% solution
           30